The demand for boxes of nails is estimated to be Q = 185 – 10 p + 3 Y, where income is measured in thousands of dollars. If p = 5, and Y = 30,

a. What is the income elasticity? Interpret and explain your result. What type of good is this?

b. How would the income elasticity change if the price were increased to $9.50? Interpret and explain your result.


1
Expert's answer
2021-11-22T10:04:57-0500

a.

income elasticity"=\frac{\%\space \Delta \space in\space Q\space demanded}{\%\space \Delta \space in\space income\space level \space Y}"

"e=\frac{3(30)}{(185-10(5)+3(30)}"

"e=\frac{90}{225}=0.4"

as income elasticity is greater than zero, the good type is normal

b.

if the price increased to $9.50

"e=\frac{3(30)}{(185-10(9.5)+3(30)}"

"e=\frac{90}{180}=0.5"

As price rises, the good becomes costly